OpenStudy (anonymous):

Which of the following are the x and y intercepts of the equation 2x - 2y = 3? a. (1.5,0), (0, -1.5) b. (2/3, 0), (0, -2/3) c. (2, 0), (0, 2) d. (3, 0), (0, -3)

OpenStudy (anonymous):

a.(1.5,0),(0,1.5)

OpenStudy (radar):

Let x=0 and solve for y then Let y=0 and solve for x. It is that easy.

OpenStudy (radar):

I will do one 2(0) -2y = 3 (Assigning 0 to x. 0-2y = 3 Y=-3/2 or -1.5 Can you do the other we know one is (0,-1.5)

OpenStudy (radar):

Compare with mine, letting y=0 2x-2(0)=3 2x=3 x=3/2=1.5 or (1.5,0)